OpenStudy (anonymous):

if a and 1/a are the zeroes of the polynomial 4x^-2x+(k-4),find the value if k.

OpenStudy (maheshmeghwal9):

\[a*1/a=k-4/4\]

OpenStudy (maheshmeghwal9):

This is product of roots & get the answer:)

OpenStudy (anonymous):

can u do it and show :P

OpenStudy (anonymous):

wot means the zeroes!?

OpenStudy (anonymous):

\[4x ^{-2x}+k-4=0\]!!!

OpenStudy (maheshmeghwal9):

very curious!!!!!!!!!!!!!

OpenStudy (anonymous):

u didn't get

OpenStudy (maheshmeghwal9):

ya I have got it but my solution is wrong since I didn't see the question clearly:\

OpenStudy (anonymous):

k